# Chemistry-Inspired Diffusion with Non-Diffe-rentiable Guidance

### v1

@@ F.2 Generalization Analysis of Neural Regressors
Figure 6: Percentage change of $\alpha$ predicted by the neural regressor when adding various noise.
@@
As observed in Table 11 in Appendix I and Table 13 in Appendix J, our results are not always stable and sometimes collapse (see the collapse analysis in Section F.4). In this section, we provide an analysis of the neural network used for gradient calculation, with the results shown in Fig. 6. We sampled 1,000 molecules from QM9 and added Gaussian noise with varying variance (1e-4, 1e-3, 1e-2, 1.0, 2.0, 3.0, 4.0) to their positions. We then evaluated the neural regressor and xTB, calculating the percentage change in predicted values from those predicted when no noise was added. For visualization purposes, we clipped the maximum percentage change to be less than 500%. The line represents the mean of the percentage changes, while the shaded area (magnitude) represents the range from maximum to minimum. As the noise scale increases, the regressors become increasingly unstable, with percentage changes reaching as high as 100%. This instability can explain the fluctuating MAE in our results, as we use the neural regressor during the generation process, where the molecules are noisier and less optimized than those in the QM9 dataset. Consequently, the regressor may provide suboptimal or even incorrect guidance during the early stages of the generation.
@@
By comparing the performance of the regressor and xTB, it is clear that xTB provides more stable predictions, as both the mean and the magnitude of instability remain lower than those of the regressor. This helps explain the performance differences in Table 4 and Table 6, where ChemGuide performs worse than $\alpha$: the reegressor produces more stable results in $\alpha$. However, ChemGuide performs better in $\Delta\epsilon$ because the regressor becomes more stable at larger noise scales, and xTB, on average, begins to “prune” invalid molecules when the noise scale is large, ceasing to provide guidance (i.e., gradients) to these invalid molecules.
@@
F.3 Generalization Analysis of Conditional GeoLDM
@@
F.4 Analysis of Gradient Collapse
@@
It’s observed in Table. 11 in Appendix I that some gradients of noisy neural guidance and clean neural guidance collapse for some scales. For example, scale 5, 20, 25, 30 when optimizing $\alpha$, so we explain why such collapses happen in this section via a case study of optimizing $\alpha$ with scale 10 (optimal scale with no collapse), 30, 40, and 50. As we increase the scale, the MAE grows quickly in the first 50 steps and the molecules collapse because the added gradient is so high that it pulls the atom apart and distorts the molecule structure. Hence, the guidance scale for each property must be carefully searched and selected so that it provides strong enough guidance while maintaining a valid molecule structure.
F.5 More Clean Guidance Optimization Steps
@@
To investigate the effect of additional recurrent steps and to prevent a single collapse from stopping the entire generation process, we divide the molecules into groups of 50 and calculate the MAE on a group-wise basis. In the figure, ”High” indicates that the MAE is too large to be displayed properly, and a circle ($\circ$) means that the MAE is either too low or too high; thus, it is classified as an outlier. For instance, in the case of $\mu$, most MAEs are high when GD step=5, except for one group with an MAE of 0.12. In terms of the lowest MAE a group can achieve, using more GD steps tends to decrease the MAE, as observed for $\alpha$ and $\Delta\epsilon$, where the lowest MAE is reduced by approximately 30% for $\alpha$ and 55% for $\mu$ (even though the lowest MAE for $\mu$ is already small, it can still be significantly reduced from 0.20 with GD step=1 to 0.09 with GD steps=3). However, in terms of average MAE, since clean guidance inherently has higher instability during generation (with details and analysis of this instability available in Fig.3 and Section I), adding more GD steps amplifies this effect. This is most evident in the case of $C_{v}$, where all groups with GD steps=3 and 10 collapse, and the MAE becomes exceptionally high when GD=5. Moreover, adding more steps does not guarantee a monotonically better MAE, as seen in $\epsilon_{LUMO}$, where 3 steps yield the optimal result while steps=5 increases the variance and mean. Therefore, selecting the number of GD steps and guidance scales carefully for each property is essential, which we leave for future work.
@@
Figure 9: Box plots of MAEs of 500 generated molecules using clean guidance with more GD steps to optimize all 6 properties. First, second (median), third quantile are plotted.
Appendix G Bilevel Optimization with Clean Guidance
@@
As shown in Section 4.3.2, our clean guidance method can achieve significantly lower MAE than the baseline. However, its ability to drastically lower MAE also contributes to unstable molecular geometries. To illustrate this, we pick two properties as a case study from Table 4 which are easiest ($\mu$) and hardest ($\alpha$) to optimize using clean guidance, and the results are shown in Table 8. Comparing the MAE and force RMS across two experiments in each property optimization, we observe that the bilevel optimization framework helps reduce force RMS and MAE at the same time, indicating that optimizing force/property yields better results on the other. However, we can observe that the force RMS of bilevel optimization is still higher than the baseline (right block of Table 8, with 200 molecules generated), but this is because the force RMS of clean guidance is much higher than the baseline (left block of Table 8, with 500 molecules generated); thus, ChemGuide in bilevel optimization can’t minimize the force significantly to beat the baseline, which is the natural hardness of optimizing force after clean guidance as the clean guidance is so strong that it significantly changes the molecular structures, making it challenging for xTB to optimize. We notice a trade-off between force and property optimization when comparing the performance across $\alpha$ and $\mu$ in bilevel optimization: lower force yields higher MAE while lower MAE brings higher force. Hence, given the difficulty of optimizing force and MAE at the same time to beat the baseline, it’s important and tricky to pick the best scales or schedules for force and property optimization, which we leave to future work. Similarly, we report full numerical results of the case study on $\alpha$ and $\mu$ optimization to Table 15 in Appendix J.
@@
Appendix H Comprehensive results on non-differentiable oracle guidance
@@
Figure 10: Histograms and distributions of force RMS and energy above ground state of 500 generated molecules from QM9 dataset using unconditional GeoLDM with ChemGuide using various scales. Energy change refers to energy above ground state.
@@
Appendix I Comprehensive results on neural guidance
@@
We present full numeric results of unconditional GeoLDM using noisy and clean neural guidance with various scales in Table 11. It is observed that larger scales yield better results while smaller scales produce results closer to unconditional GeoLDM. This is intuitive because larger scales provide more guidance to the generation, which brings better properties; on the other hand, since we are using unconditional GeoLDM as the backbone, the guidance becomes more negligible when the scale is small and thus the performance converges to unconditional GeoLDM. ✗  in Table 11 means that the molecules collapse in the middle of the generation, so there are no output molecules. Explanation and case study of collapse can be found in Section F.4 and Section F.2. To better investigate the effect of larger scales, we split 500 generated molecules into 10 groups so that a single collapse will not stop the generation of all molecules, we put the remaining number of molecules in the parenthesis in Table 11 and report the MAE of the molecules in groups without collapse. For example, 3.2309 (450) in $\alpha$ of scale=5 means there are 450 molecules in groups with no collapse and their MAE is 3.2309.
@@
We plot the MAE trajectories of optimizing all six properties in Fig. 11. Applying noisy neural guidance results in much smoother trajectories compared to the baseline. It reduces MAE significantly during the first 50 epochs and then stabilizes, making noisy guidance most effective in the early stages of generation. Clean neural guidance also lowers MAE in the first 100 steps to some extent, but it fluctuates more than noisy guidance (and sometimes more than the baseline, as in the trajectory of optimizing $\epsilon_{LUMO}$). However, clean guidance becomes more powerful in the final 50 steps, where we observe a sharp drop in MAE across all six properties. The difference in smoothness is because, in noisy guidance, the gradient is propagated through the denoising estimation (Eq.8) and the VAE decoder (Algorithm1), inherently accounting for the effect of noise in the gradient calculation, which leads to consistent guidance throughout the generation steps. On the contrary, in clean guidance, the gradient is computed in the clean space (Eq.11) and then projected back into the generation process by adding noise (Eq.10). Since the gradient does not propagate through the denoising estimation and VAE decoder, it is more sensitive to noise, resulting in greater fluctuations and inconsistency during guidance. Noisy guidance stabilizes and doesn’t exhibit a sharp drop in later steps because the loss decreases as the MAE reduces in the early stages, so the gradient is small, which mainly serves to maintain the current molecular structure, preventing the MAE from converging to that of GeoLDM. This can be observed in Table 11, where the MAE for smaller scales is closer to that of unconditional GeoLDM. Although clean guidance helps reduce MAE in earlier steps of generation, it reaches its full potential in the later steps, when the molecules are more similar and close to decoded molecules, leading to the sharp drop in MAE.
@@
Appendix J Comprehensive results on bilevel optimization
@@
The MAE and force metrics are shown in Table 13 and Table 12 respectively. Since we only add ChemGuide to C-GeoLDM, our expectation is that we can achieve lower forces without increasing MAE. However, the results go beyond our expectations: our method not only achieves better force RMS, energy above ground state, and validity, but also reduces MAE in the first 5 properties. This shows the effectiveness of combining explicit property optimization with ChemGuide, and verifies the feasibility of our bilevel optimization framework.
@@
To illustrate how bilevel optimization works, we provide MAE trajectories of each property in Fig. 12. Compared to MAE trajectories in unconditional GeoLDM with noisy guidance in Fig. 11, the trajectories of bilevel optimization with noisy guidance behave similarly in the sense that the MAE is reduced significantly during the first 50 guidance steps and stabilizes. The difference is that the trajectories of bilevel optimization with noisy guidance have slightly larger fluctuation during the early steps of guidance. We reason that in the early steps, the molecules are less stable, so ChemGuide provides larger guidance such that the guidance from ChemGuide and noisy guidance interact with each other and result in fluctuated trajectories. The trajectories are also more converged to GeoLDM because the generation process leans less toward a lower MAE, as it keeps a balance between property and stability optimization, which can be verified from 14.
@@
Finally, we present the algorithm of bilevel optimization with clean guidance and ChemGuide in Algorithm 2. The detailed results of optimizing $\alpha$ and $\mu$ are shown in Table 15. Recall that $\alpha$ is the most challenging to optimize and $\mu$ is the easiest according to Table 4. In both properties, it’s obvious that increasing guidance scales results in lower MAE yet larger force RMS, which is because the effect of clean guidance exceeds that of ChemGuide, so it’s a trade-off between optimizing force and property. Comparing the two methods of the two properties, we notice that optimizing $\alpha$ yields better force RMS yet higher MAE while optimizing $\mu$ results in worse force RMS yet lower MAE, which is another trade-off between optimizing force and property. In each property, although we can’t directly compare the performance of GeoLDM with clean guidance and bilevel optimization with clean guidance because we sample 500 molecules from the former and 200 from the latter, we can compare them with their respective baselines. We notice that by adding ChemGuide on force in bilevel optimization, we achieve lower force RMS percentage changes from the baselines. Surprisingly, the MAE of bilevel optimization with clean guidance is lower than that of GeoLDM with clean guidance, which seems to conflict with our findings in bilevel optimization with noisy guidance, we suspect that this is because, as the molecules are optimized to be more stable and thus more similar to the denoised ones, clean guidance can obtain more power in reducing the MAE, as explained in Section I. However, there is a collapse in the bilevel optimization of $\mu$ with scale 10 while it doesn’t happen in GeoLDM with clean guidance, we suspect that this is due to the exceptionally high force RMS in GeoLDM with clean guidance, so ChemGuide provides much higher gradient and the guidance from ChemGuide (pushing toward lower force RMS) and clean guidance (pushing towards higher force RMS) conflict, resulting in the collapse of molecular structures. Nevertheless, although the force RMS of bilevel optimization with clean guidance is still higher than the baselines, it is already high in GeoLDM with clean guidance and we can reduce it by adding ChemGuide, which, together with lower MAE, indicates the feasibility of combining ChemGuide with clean guidance and our success in obtaining both better property and force metrics.
@@
Energy above ground state (Eh) -0.0083 -0.0085 -0.0079 (-13.93$\downarrow$) -0.0084 -0.0104 -0.0094 -0.0092
@@
Energy above ground state (Eh) -0.0106 (16.72$\uparrow$) -0.0106 -0.0112 -0.0111 -0.0116 -0.0113 -0.0091
@@
Energy above ground state (Eh) -0.0067 -0.0066 -0.0063 (-21.39$\downarrow$) -0.0066 -0.0092 -0.0097 -0.0081
@@
Energy above ground state (Eh) -0.0057 (-27.87$\downarrow$) -0.0057 -0.0061 -0.0060 -0.0125 -0.0090 -0.0079
@@
Energy above ground state (Eh) -0.0079 (-0.60$\downarrow$) -0.0079 -0.0079 -0.0102 -0.0096 -0.0083 -0.0079
@@
Energy above ground state (Eh) -0.0061 (-16.41$\downarrow$) -0.0065 -0.0064 -0.0065 -0.0094 -0.0072 -0.0073
@@
Table 14: Force Metrics of all 6 properties of 200 sampled molecules from implicit bilevel optimization with noisy guidance. The ”Top” scales of each property can be found in Table. 11. We use C-EDM and C-GeoLDM as baselines. * and bold denote the overall best and the best within different scales, respectively. Percentage changes between our results and C-GeoLDM are shown in parenthesis
@@
Energy above ground state (Eh) -0.0087 -0.0047 -0.0043 -0.0028 (-69.26$\downarrow$) -0.0052 (-43.72$\downarrow$) -0.0094 -0.0092
@@
Energy above ground state (Eh) -0.0035 -0.0108 -0.0027 -0.0025 (-72.31$\downarrow$) -0.0050 (-45.45$\downarrow$) -0.0113 -0.0091
@@
Energy above ground state (Eh) -0.0139 -0.0036 -0.0037 -0.0025 (-68.78$\downarrow$) -0.0060 (-24.96$\downarrow$) -0.0097 -0.0081
@@
Energy above ground state (Eh) -0.0032 -0.0028 -0.0030 -0.0028 (-64.91$\downarrow$) -0.0029 (-62.59$\downarrow$) -0.0090 -0.0079
@@
Energy above ground state (Eh) -0.0021 (-73.79$\downarrow$) -0.0024 -0.0026 -0.0025 -0.0024 (-69.62$\downarrow$) -0.0083 -0.0079

### v2

@@ H.4 Generalization Analysis of Neural Regressors
Figure 7: Percentage change of $\alpha,\mu,\epsilon_{\text{HOMO}},\epsilon_{\text{LUMO}},\Delta_{\epsilon}$ predicted by the neural regressor when adding various noise.
@@
As observed in Table 16 in Appendix L and Table 18 in Appendix M, our results are not always stable and sometimes collapse (see the collapse analysis in Section H.6). In this section, we provide an analysis of the neural network used for gradient calculation, with the results shown in Fig. 7. We sampled 1,000 molecules from QM9 and added Gaussian noise with varying variances (1e-4, 1e-3, 1e-2, 1.0, 2.0, 3.0, 4.0) to their positions. We then evaluated the neural regressor and xTB, calculating the percentage change in predicted values from those predicted when no noise was added. For visualization purposes, we clipped the maximum percentage change to be less than 500%. The line represents the mean of the percentage changes, while the shaded area (magnitude) represents the range from maximum to minimum. As the noise scale increases, the regressors become increasingly unstable, with percentage changes reaching as high as 100%. This instability can explain the fluctuating MAE in our results, as we use the neural regressor during the generation process, where the molecules are noisier and less optimized than those in the QM9 dataset. Consequently, the regressor may provide suboptimal or even incorrect guidance during the early stages of the generation.
@@
By comparing the performance of the regressor and xTB, it is clear that xTB provides more stable predictions, as both the mean and the magnitude of instability remain lower than those of the regressor. This helps explain the performance differences in Table 4 and Table 6, where ChemGuide performs worse than $\alpha$: the regressor produces more stable results in $\alpha$. However, ChemGuide performs better in $\Delta\epsilon$ because the regressor becomes more stable at larger noise scales, and xTB, on average, begins to “prune” invalid molecules when the noise scale is large, ceasing to provide guidance (i.e., gradients) to these invalid molecules.
@@
H.5 Generalization Analysis of Conditional GeoLDM
@@
H.6 Analysis of Gradient Collapse
@@
It’s observed in Table. 16 in Appendix L that some gradients of noisy neural guidance and clean neural guidance collapse for some scales. For example, scale 5, 20, 25, 30 when optimizing $\alpha$, so we explain why such collapses happen in this section via a case study of optimizing $\alpha$ with scale 10 (optimal scale with no collapse), 30, 40, and 50. As we increase the scale, the MAE grows quickly in the first 50 steps and the molecules collapse because the added gradient is so high that it pulls the atom apart and distorts the molecule structure. Hence, the guidance scale for each property must be carefully searched and selected so that it provides strong enough guidance while maintaining a valid molecule structure.
H.7 More Clean Guidance Optimization Steps
@@
To investigate the effect of additional recurrent steps and to prevent a single collapse from stopping the entire generation process, we divide the molecules into groups of 50 and calculate the MAE on a group-wise basis. In the figure, ”High” indicates that the MAE is too large to be displayed properly, and a circle ($\circ$) means that the MAE is either too low or too high; thus, it is classified as an outlier. For instance, in the case of $\mu$, most MAEs are high when GD step=5, except for one group with an MAE of 0.12. In terms of the lowest MAE a group can achieve, using more GD steps tends to decrease the MAE, as observed for $\alpha$ and $\Delta\epsilon$, where the lowest MAE is reduced by approximately 30% for $\alpha$ and 55% for $\mu$ (even though the lowest MAE for $\mu$ is already small, it can still be significantly reduced from 0.20 with GD step=1 to 0.09 with GD steps=3). However, in terms of average MAE, since clean guidance inherently has higher instability during generation (with details and analysis of this instability available in Fig.3 and Section L), adding more GD steps amplifies this effect. This is most evident in the case of $C_{v}$, where all groups with GD steps=3 and 10 collapses, and the MAE becomes exceptionally high when GD=5. Moreover, adding more steps does not guarantee a monotonically better MAE, as seen in $\epsilon_{LUMO}$, where 3 steps yield the optimal result while steps=5 increase the variance and mean. Therefore, selecting the number of GD steps and guidance scales carefully for each property is essential, which we leave for future work.
@@
Figure 10: Box plots of MAEs of 500 generated molecules using clean guidance with more GD steps to optimize all 6 properties. The first, second (median), and third quantiles are plotted.
Appendix I Other Optimization Algorithms
In this section, we compare our proposed ChemGuide with an alternative class of gradient-free methods that incorporate non-differentiable oracles into the diffusion process: evolutionary algorithms (Schneuing et al., 2022; Huang et al., 2024). The evolutionary algorithm, detailed in Algorithm 2, has two main parameters: the variant size ($k$) and the evolution interval ($E$). During the diffusion process, the population is preserved and evolution is performed at fixed intervals. At each evolution step, $k-1$ noise perturbations are added to the population, resulting in $k$ variants for each molecule, where the unperturbed molecule is treated as a variant. The best variant is then selected as the new population based on evaluations from the non-differentiable oracle (i.e. xTB). Specifically, the oracle calculates the force RMS, and the variant with the lowest force RMS is selected.
We explored several combinations of variant size and evolution interval on QM9, and the results are summarized in Table 12. For ease of comparison, the results using the best scales of ChemGuide are also provided in Table 1. We can observe ChemGuide is on par with the evolutionary algorithm in validity and stability, and significantly outperforms it in terms of force RMS and energy above the ground state. This advantage can be attributed to the fact that gradients provide informative guidance by indicating the correct direction for optimizing stability. In contrast, the evolutionary algorithm operates without relying on gradients, which results in a lack of directional guidance. Consequently, our approach enables controllable optimization with this gradient-based guidance. While the evolutionary algorithm selects the best variant at each evolutionary step, its optimization direction is essentially random, making the process less controllable. Additionally, the gradient inherently adjusts the direction of the generative process, whereas the evolutionary algorithm does not dynamically adapt its direction. Therefore, our ChemGuide offers a more generic and inherently guided optimization process.
Algorithm 2 ChemGuide diffusion sampling with evolutionary algorithm
  Input: a latent diffusion model $\epsilon_{\theta}$, a VAE decoder $\mathcal{D}$, a composition function $\mathcal{F}$, target property score $y$, variant size $k$, variant scale $s_{v}$, evolution interval $E$.
  Output: optimized molecule $[\mathbf{x},\mathbf{h}]$
  ${\bm{z}}_{T}\leftarrow\mathcal{N}(0,\mathbf{I})$
  $population\leftarrow[{\bm{z}}_{t}]$
  for all $t$ from $T$ to 1 do
     // Add variants
     if $t\,\%E=0$ then
        $population$.extend$([{\bm{z}}_{t}+s_{v}\cdot\epsilon_{1},{\bm{z}}_{t}+s_{v}\cdot\epsilon_{2},\cdots,{\bm{z}}_{t}+s_{v}\cdot\epsilon_{k-1}])$, where $\epsilon_{1},\cdots,\epsilon_{k-1}\sim\mathcal{N}(0,\mathbf{I})$
     end if
     for all ${\bm{z}}_{t,i}\in population$ do
        $\mu_{t-1,i},\Sigma_{t-1,i}\leftarrow\frac{1}{\sqrt{1-\beta_{t}}}({\bm{z}}_{t,i}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}^{2}}}\epsilon_{\theta}({\bm{z}}_{t,i},t)),\sigma_{t}^{2}\mathbf{I}$ (Eq. 3)
        ${\bm{z}}_{t-1,i}\leftarrow\mathcal{N}(\mu_{t-1,i},\Sigma_{t-1,i})$ (Eq. 6)
        Replace ${\bm{z}}_{t,i}$ in $population$ by ${\bm{z}}_{t-1,i}$
     end for
     ${\bm{z}}_{t-1}^{*}\leftarrow\operatorname*{arg\,min}_{{\bm{z}}_{t-1}\in population}\|\mathcal{F}({\bm{z}}_{t-1})\|^{2}$ // Select the best variant
     $population\leftarrow[{\bm{z}}_{t-1}^{*}]$
  end for
  $[\mathbf{x},\mathbf{h}]\leftarrow\mathcal{D}({\bm{z}}_{0}^{*})$
  return  $[\mathbf{x},\mathbf{h}]$
Table 12: Metrics of 500 generated molecules from QM9 using GeoLDM with evolutionary algorithm (Algorithm 2). $k,E$ denote the variant size and evolution interval. * and bold denote the overall best result and our best result, respectively. Percentage changes between our results and GeoLDM are shown in parentheses. For ease of comparison, we report the results using the best scale of ChemGuide from Table 1.
Metric (variant size, evolution interval) ChemGuide Baseline
($k=3$, $E=20$) ($k=3$, $E=50$) ($k=5$, $E=20$) ($k=5$, $E=50$) scale = 0.0001 EDM GeoLDM
Force RMS (Eh/Bohr) 0.0112 0.0111 0.0110 0.0109 (-1.75$\downarrow$) 0.0104 (-6.76$\downarrow$) 0.0114 0.0111
Validity 91.60% 87.60% 90.00% 92.40% (2.60$\uparrow$) 91.40% (1.60$\uparrow$) 86.60% 89.80%
Uniqueness 100.00% 100.00% 99.78% 100.00% 100.00% 100.00% 100.00%
Atom Stability 98.95% 98.76% 98.66% 99.14% (0.21$\uparrow$) 99.02% (0.09$\uparrow$) 98.53% 98.93%
Molecule Stability 90.00% 87.60% 88.00% 91.20% (2.40$\uparrow$) 90.60% (1.80$\uparrow$) 83.60% 88.80%
Energy above ground state (Eh) 0.0048 (-4.45$\downarrow$) 0.0053 0.0056 0.0054 0.0042 (-15.78$\downarrow$) 0.0072 0.0050
Appendix J Bilevel Optimization with Clean Guidance
@@ Appendix I Other Optimization Algorithms
As shown in Section 4.3.2, our clean guidance method can achieve significantly lower MAE than the baseline. However, its ability to drastically lower MAE also contributes to unstable molecular geometries. To illustrate this, we pick two properties as a case study from Table 4 which are easiest ($\mu$) and hardest ($\alpha$) to optimize using clean guidance, and the results are shown in Table 13. Comparing the MAE and force RMS across two experiments in each property optimization, we observe that the bilevel optimization framework helps reduce force RMS and MAE at the same time, indicating that optimizing force/property yields better results on the other. However, we can observe that the force RMS of bilevel optimization is still higher than the baseline (right block of Table 13, with 200 molecules generated), but this is because the force RMS of clean guidance is much higher than the baseline (left block of Table 13, with 500 molecules generated); thus, ChemGuide in bilevel optimization can’t minimize the force significantly to beat the baseline, which is the natural hardness of optimizing force after clean guidance as the clean guidance is so strong that it significantly changes the molecular structures, making it challenging for xTB to optimize. We notice a trade-off between force and property optimization when comparing the performance across $\alpha$ and $\mu$ in bilevel optimization: lower force yields higher MAE while lower MAE brings higher force. Hence, given the difficulty of optimizing force and MAE at the same time to beat the baseline, it’s important and tricky to pick the best scales or schedules for force and property optimization, which we leave to future work. Similarly, we report full numerical results of the case study on $\alpha$ and $\mu$ optimization to Table 20 in Appendix M.
@@
Appendix K Comprehensive Results on Non-differentiable Oracle Guidance
@@
Figure 11: Histograms and distributions of force RMS and energy above the ground state of 500 generated molecules from QM9 dataset using unconditional GeoLDM with ChemGuide using various scales. Energy change refers to energy above ground state.
@@
Appendix L Comprehensive Results on Neural Guidance
@@
We present full numeric results of unconditional GeoLDM using noisy and clean neural guidance with various scales in Table 16. It is observed that larger scales yield better results while smaller scales produce results closer to unconditional GeoLDM. This is intuitive because larger scales provide more guidance to the generation, which brings better properties; on the other hand, since we are using unconditional GeoLDM as the backbone, the guidance becomes more negligible when the scale is small and thus the performance converges to unconditional GeoLDM. ✗  in Table 16 means that the molecules collapse in the middle of the generation, so there are no output molecules. Explanation and case study of collapse can be found in Section H.6 and Section H.4. To better investigate the effect of larger scales, we split 500 generated molecules into 10 groups so that a single collapse will not stop the generation of all molecules, we put the remaining number of molecules in the parenthesis in Table 16 and report the MAE of the molecules in groups without collapse. For example, 3.2309 (450) in $\alpha$ of scale=5 means there are 450 molecules in groups with no collapse and their MAE is 3.2309.
@@
We plot the MAE trajectories of optimizing all six properties in Fig. 12. Applying noisy neural guidance results in much smoother trajectories compared to the baseline. It reduces MAE significantly during the first 50 epochs and then stabilizes, making noisy guidance most effective in the early stages of generation. Clean neural guidance also lowers MAE in the first 100 steps to some extent, but it fluctuates more than noisy guidance (and sometimes more than the baseline, as in the trajectory of optimizing $\epsilon_{LUMO}$). However, clean guidance becomes more powerful in the final 50 steps, where we observe a sharp drop in MAE across all six properties. The difference in smoothness is because, in noisy guidance, the gradient is propagated through the denoising estimation (Eq.8) and the VAE decoder (Algorithm2), inherently accounting for the effect of noise in the gradient calculation, which leads to consistent guidance throughout the generation steps. On the contrary, in clean guidance, the gradient is computed in the clean space (Eq.11) and then projected back into the generation process by adding noise (Eq.10). Since the gradient does not propagate through the denoising estimation and VAE decoder, it is more sensitive to noise, resulting in greater fluctuations and inconsistency during guidance. Noisy guidance stabilizes and doesn’t exhibit a sharp drop in later steps because the loss decreases as the MAE reduces in the early stages, so the gradient is small, which mainly serves to maintain the current molecular structure, preventing the MAE from converging to that of GeoLDM. This can be observed in Table 16, where the MAE for smaller scales is closer to that of unconditional GeoLDM. Although clean guidance helps reduce MAE in earlier steps of generation, it reaches its full potential in the later steps, when the molecules are more similar and close to decoded molecules, leading to a sharp drop in MAE.
@@
Appendix M Comprehensive Results on Bilevel Optimization
@@
The MAE and force metrics are shown in Table 18 and Table 17 respectively. Since we only add ChemGuide to C-GeoLDM, our expectation is that we can achieve lower forces without increasing MAE. However, the results go beyond our expectations: our method not only achieves better force RMS, energy above ground state, and validity but also reduces MAE in the first 5 properties. This shows the effectiveness of combining explicit property optimization with ChemGuide, and verifies the feasibility of our bilevel optimization framework.
@@
To illustrate how bilevel optimization works, we provide MAE trajectories of each property in Fig. 13. Compared to MAE trajectories in unconditional GeoLDM with noisy guidance in Fig. 12, the trajectories of bilevel optimization with noisy guidance behave similarly in the sense that the MAE is reduced significantly during the first 50 guidance steps and stabilizes. The difference is that the trajectories of bilevel optimization with noisy guidance have slightly larger fluctuation during the early steps of guidance. We reason that in the early steps, the molecules are less stable, so ChemGuide provides larger guidance such that the guidance from ChemGuide and noisy guidance interact with each other and result in fluctuated trajectories. The trajectories are also more converged to GeoLDM because the generation process leans less toward a lower MAE, as it keeps a balance between property and stability optimization, which can be verified from 19.
@@
Finally, we present the algorithm of bilevel optimization with clean guidance and ChemGuide in Algorithm 3. The detailed results of optimizing $\alpha$ and $\mu$ are shown in Table 20. Recall that $\alpha$ is the most challenging to optimize and $\mu$ is the easiest according to Table 4. In both properties, it’s obvious that increasing guidance scales results in lower MAE yet larger force RMS, which is because the effect of clean guidance exceeds that of ChemGuide, so it’s a trade-off between optimizing force and property. Comparing the two methods of the two properties, we notice that optimizing $\alpha$ yields better force RMS yet higher MAE while optimizing $\mu$ results in worse force RMS yet lower MAE, which is another trade-off between optimizing force and property. In each property, although we can’t directly compare the performance of GeoLDM with clean guidance and bilevel optimization with clean guidance because we sample 500 molecules from the former and 200 from the latter, we can compare them with their respective baselines. We notice that by adding ChemGuide on force in bilevel optimization, we achieve lower force RMS percentage changes from the baselines. Surprisingly, the MAE of bilevel optimization with clean guidance is lower than that of GeoLDM with clean guidance, which seems to conflict with our findings in bilevel optimization with noisy guidance, we suspect that this is because, as the molecules are optimized to be more stable and thus more similar to the denoised ones, clean guidance can obtain more power in reducing the MAE, as explained in Section L. However, there is a collapse in the bilevel optimization of $\mu$ with scale 10 while it doesn’t happen in GeoLDM with clean guidance, we suspect that this is due to the exceptionally high force RMS in GeoLDM with clean guidance, so ChemGuide provides much higher gradient and the guidance from ChemGuide (pushing toward lower force RMS) and clean guidance (pushing towards higher force RMS) conflict, resulting in the collapse of molecular structures. Nevertheless, although the force RMS of bilevel optimization with clean guidance is still higher than the baselines, it is already high in GeoLDM with clean guidance and we can reduce it by adding ChemGuide, which, together with lower MAE, indicates the feasibility of combining ChemGuide with clean guidance and our success in obtaining both better property and force metrics.
@@
Energy above ground state (Eh) 0.0083 0.0085 0.0079 (-13.93$\downarrow$) 0.0084 0.0104 0.0094 0.0092
@@
Energy above ground state (Eh) 0.0106 (16.72$\uparrow$) 0.0106 0.0112 0.0111 0.0116 0.0113 0.0091
@@
Energy above ground state (Eh) 0.0067 0.0066 0.0063 (-21.39$\downarrow$) 0.0066 0.0092 0.0097 0.0081
@@
Energy above ground state (Eh) 0.0057 (-27.87$\downarrow$) 0.0057 0.0061 -0.0060 0.0125 0.0090 0.0079
@@
Energy above ground state (Eh) 0.0079 (-0.60$\downarrow$) 0.0079 0.0079 0.0102 0.0096 0.0083 0.0079
@@
Energy above ground state (Eh) 0.0061 (-16.41$\downarrow$) 0.0065 0.0064 -0.0065 0.0094 0.0072 0.0073
@@
Table 19: Force Metrics of all 6 properties of 200 sampled molecules from implicit bilevel optimization with noisy guidance. The “Top” scales of each property can be found in Table. 16. We use C-EDM and C-GeoLDM as baselines. * and bold denote the overall best and the best within different scales, respectively. Percentage changes between our results and C-GeoLDM are shown in parenthesis
@@
Energy above ground state (Eh) 0.0087 0.0047 0.0043 0.0028 (-69.26$\downarrow$) 0.0052 (-43.72$\downarrow$) 0.0094 0.0092
@@
Energy above ground state (Eh) 0.0035 0.0108 0.0027 0.0025 (-72.31$\downarrow$) 0.0050 (-45.45$\downarrow$) 0.0113 0.0091
@@
Energy above ground state (Eh) 0.0139 0.0036 0.0037 0.0025 (-68.78$\downarrow$) 0.0060 (-24.96$\downarrow$) 0.0097 0.0081
@@
Energy above ground state (Eh) 0.0032 0.0028 0.0030 0.0028 (-64.91$\downarrow$) 0.0029 (-62.59$\downarrow$) 0.0090 0.0079
@@
Energy above ground state (Eh) 0.0021 (-73.79$\downarrow$) 0.0024 0.0026 0.0025 0.0024 (-69.62$\downarrow$) 0.0083 0.0079